Circular No. 5804

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET HOLMES (1993i) S. Nakano, Sumoto, Japan, reports the recovery of this comet by T.

Seki at Geisei, the positions being very close to the prediction on MPC 16379 (ephemeris on MPC 21958). The comet is diffuse with little central condensation, coma diameter 20′′.

1993 UT  α_{2000}  δ_{2000}  m_{1}
May 24.77222  1^{h}12^{m}09.^{s}20  +17^{o}34′18′′.7  18
25.77639  1 13 59.40  +17 51 18.8  18

HD 98800 F. C. Fekel, Marshall Space Flight Center, NASA; and B. W. Bopp,

University of Toledo, write: “High-resolution CCD spectra of the dusty K5 V star HD 98800 (Zuckerman and Becklin 1993, Ap.J. 406, L33; Gregorio-Hetem et al. 1992, A.J. 103, 549) have been obtained at Kitt Peak National Observatory. The spectra confirm the dwarf spectral type and show rotationally-broadened photospheric lines, Hα emission, and a very strong Li I absorption line. The overall spectrum of HD 98800 closely resembles that of the young, chromospherically active, single BY-Dra star HD 82558 (Fekel et al. 1986, A.J. 92, 1150). The measured Li I equivalent width of 0.04 nm implies an abundance log n(Li) = 2.2 and an age of 20-30 Myr. An upper limit on the rotation period, based on our measured v sin i = 11 ± 1 km/s, is 3.7 days. HD 98800 should be photometrically variable, and potentially a strong x-ray and radio source.”

1993 KA E. S. Howell and M. C. Nolan, Lunar and Planetary Laboratory, com-

municate: “We obtained CCD photometry of 1993 KA using the 1.54-m Catalina telescope. We measure V = 16.916 ± 0.014 and V–R = +0.434 ± 0.018, averaged over the interval May 18.2125-18.2800 UT. In this time interval, the brightness variation in both Harris V and Cousins R filters was < 0.10 mag, which is within 1.5σ of the mean value. However, the Harris B filter photometry shows significant variation in this same interval; B–V ranges from +0.482 ± 0.046 to +0.747 ± 0.030 in ∼ 30 min. We did not obtain data over a full rotation period, and we can only estimate that the period is longer than 3 hr. CCD spectra were obtained on May 19 and 20, which may help in understanding these color variations.”

1993 May 26 Daniel W. E. Green